# On action rate admissibility criteria

### v3

@@ 1 Introduction
The paper [9] introduces a version of the least action principle as a selection criterion in initial value problems possessing non-unique solutions. Interest arises from examples that use convex integration techniques to obtain families of non-unique wild solutions to the Euler system (De Lellis and Székelyhidi [4, 5]), and non-unique solutions to the Navier-Stokes equations (Buckmaster and Vicol [1]). Convex integration is also employed by Chiodaroli and Kreml [2] to investigate the two-dimensional Riemann initial value problem for the isentropic compressible Euler system. They construct a large set of entropic wild solutions to the Riemann problem.
In the present paper we continue the discussion of admissibility criteria started in [9] in view of the recent results of Markfelder and Pellhammer [14]. Admissibility criteria in fluid dynamics include the principles of minimum potential energy or entropy, various rate of change criteria, and that of least action (LAAP) developed in [9]. The latter principle applied to the initial value problem considered in [2] selects the 2-shock solution provided the solutions obtained by convex integration are the only other ones admitted for comparison. On the other hand, Chiodaroli and Kreml [2] show that there exist Riemann data for which there are wild solutions that possess entropy rate lower than the 2-shock solution which therefore is not selected by Dafermos’s entropy rate criterion [3]. On the other hand, in particular cases, LAAP selects the 2-shock in preference to wild solutions.
LAAP, however, does not select the 2-shock solution when the set of comparison solutions is enlarged to include other weak solutions in addition to the 2-shock and convex integration solutions. Markfelder and Pellhammer [14] recently constructed hybrid solutions to the same Riemann initial value problem for which, subject to prescribed initial data, LAAP does not identify the 2-shock solutions as being uniquely admissible. In fact, the example indicates that LAAP does not provide an admissible solution to the Riemann problem. Consequently, when the 2-shock solution is preferred in this Riemann problem, an alternative to LAAP must be sought.
@@
Section 2 collects basic definitions, states the least action admissibility principle (LAAP) derived in [9] and formulates the new least action admissibility principle (LAAP0) and two closely related rate criteria. Section 3, after a brief description of convex integration, introduces the two-dimensional Riemann initial value problem for the barotropic and isentropic compressible Euler equations of gas dynamics and summarises further material from [2] including the construction of sub-solutions. The 2-shock solution is also described. Justification for the introduction of LAAP0 is provided in Section 4, while Section 5 defines the relevant notion of action and applies LAAP0 to the Riemann problem treated in [2] to show that the 2-shock solution is admissible.
@@ 2 Least action admissibility criteria
$\frac{d{\boldsymbol{u}}}{dt}=F({\boldsymbol{u}}),\qquad{\boldsymbol{u}}(t_{0})={\boldsymbol{u}}_{0},\qquad t\in[t_{0},\,T],$ (2.1)
where $F({\boldsymbol{u}})$ is a given operator and $[t_{0},\,T]$ is the maximal time interval of existence of solutions belonging to $S$. Global existence is not assumed so that possibly $T<\infty.$ The system (2.1) is taken to be the Euler-Lagrange equation derived from a least action principle. Suppose that $S$ is not a singleton and in consequence (2.1) has non-unique solutions. The objective is to formulate an admissibility criterion defined in terms of the action that selects (not necessarily unique) elements of $S$. The specific form of action we will use is postponed to Section 5. For present purposes, it is sufficient to assume that for each solution ${\boldsymbol{u}}\in S$, the action $A({\boldsymbol{u}})(t_{0},\,t)$ is a non-decreasing function of time $t$ that vanishes initially for all ${\boldsymbol{u}}\in S$.
Further definitions are required, but before their statement, it is convenient to recall the Least Action Admissibility Principle (LAAP) introduced in [9] and applied to the Riemann initial value problem for compressible Euler systems and for Dafermos’s nonlinear oscillator [8]. For fixed final time $t_{1}$, we have:
@@
In a time interval $[t_{0},t_{1}]$, where $t_{1}\in{\mathbb{R}}$ is less than or equal to the infimum of the maximal time of existence for solutions in $S$, a solution ${\boldsymbol{u}}\in S$ is LAAP-admissible if the action $A({\boldsymbol{u}})(t_{0},t_{1})$ is not greater than the action of all other solutions in $S$. The solution ${\boldsymbol{u}}\in S$ is strictly LAAP-admissible when the inequality is strict.
In the context of the wild solutions of the Riemann problem of [2], in a number of cases LAAP identifies the 2-shock solution as the only strictly LAAP-admissible solution for all $t_{1}>t_{0}$. But if the set $S$ includes solutions other than the wild solutions constructed in [2] and the 2-shock solution, LAAP, as just explained, can be ineffective. Obviously, the action depends upon the value chosen for $t_{1}$. In fact, as discussed in the next Section, this is the main feature in the construction of Markfelder and Pellhammer [14]: some choices of $t_{1}$ allow the one-dimensional 2-shock solution to be LAAP-admissible, while others do not.
Here we formulate a local in time version of LAAP, called LAAP0, for which the 2-shock solution is admissible in the case of Riemann problems treated in [2], when the set $S$ also includes wild solutions and the hybrid solutions of [14]. To define LAAP0 we need the following definitions. (Note that in Definition 2.2 the time interval $t_{1}({\boldsymbol{v}})$ depends on the solution ${\boldsymbol{v}}$.)
@@
2. For general solutions, a definition of entropy rate is presented in [7, Sects 1.2,1.3] when the time derivative only exists almost everywhere. The corresponding general definition of the action rate is obtained by replacing the energy $E$ by the action $A$ in the discussion of [7].
@@
We apply the principles defined in Propositions 2.2 and 2.3 to the two-dimensional Riemann problem for compressible isentropic Euler equation studied in e.g., [2, 9, 14].
3 The Riemann problem
This Section defines the Riemann problem to we wish to analyse and. for convenience, summarises properties of various solutions to be studied subsequently. First, however, elements of convex integration are briefly reviewed.
@@ 3.1 Convex integration
The two main elements of convex integration pertinent to the present study comprise the concepts of a sub-solution and corrugation. For equation (2.1) a sub-solution is the function ${\boldsymbol{w}}$ that in analogy with a sub-harmonic function satisfies the inequality
@@
Corrugation is a method of successively superposing on the sub-solution a sequence of oscillations at diminishing scale that in the limit converges to an appropriately defined (weak) exact solution to the equation of interest in the sense of distributions. In Nash’s historic paper [15] on isometric embedding of Riemannian manifolds, sub-solutions correspond to short embeddings. The challenge was to produce the corrugation algorithm which he termed stages. Gromov [11, 12] developed the technique while De Lellis and Székelyhidi [4, 5] applied convex integration to construct wild solutions to the Euler equations. The lecture notes by Markfelder [13] may be consulted for details. An overview is provided in the survey by De Lellis and Székelyhidi [6].
Chiodaroli and Kreml [2] apply the convex integration procedure to the 2-dimensional Riemann problem we are dealing with and, specifically, construct relevant sub-solutions. Hence we need to explain their construction to make it easier to understand our application of LAAP0 to the isentropic form of two dimensional barotropic Euler system discussed in [2].
@@ 3.2 Isentropic compressible Euler system
In what follows, the independent variables are time $t$ and position ${\boldsymbol{x}}$ such that $(t,{\boldsymbol{x}})\in[0,\infty)\times{\mathbb{R}}^{2}$. Let $\rho$ be the unknown mass density, ${\boldsymbol{m}}$ the momentum, and $p(\rho)$ the constitutively determined pressure. The fluid velocity is ${\boldsymbol{v}}={\boldsymbol{m}}/\rho.$ In an obvious notation, the two dimensional compressible barotropic Euler system of gas dynamics is given by
$\partial_{t}\rho+div_{x}{\boldsymbol{m}}$ $=$ $0,$ (3.1)
$\partial_{t}{\boldsymbol{m}}+div_{x}\left(\frac{{\boldsymbol{m}}\otimes{\boldsymbol{m}}}{\rho}\right)+\nabla_{x}p(\rho)$ $=$ $0,$ (3.2)
to which are adjoined initial conditions for $(\rho,{\boldsymbol{v}})$:
$(\rho,\,{\boldsymbol{v}})(0,\,.)=(\rho_{0},\,{\boldsymbol{v}}_{0}).$ (3.3)
The pair $(\rho,\,{\boldsymbol{v}})$ is a weak solution to the initial value problem for the barotropic compressible Euler system when the pair satisfy (3.1)-(3.3) in the sense of distributions.
The set $S$ of solutions to the initial value problem (3.1)-(3.3) consists of all entropic weak solutions, that is, weak solutions that satisfy the energy-entropy condition in the sense of distributions (see for example [2, 3, 13, 14]):
@@
Chiodaroli and Kreml [2] study the isentropic form of the barotropic system (3.1) and (3.2) for which $p(\rho)=\rho^{\gamma}$ with constant $\gamma\geq 1$. The Riemann initial data is
$(\rho_{0},\,{\boldsymbol{m}}_{0})=\left\{\begin{array}[]{lll}(\rho_{-},\,{\boldsymbol{m}}_{-}),&x_{2}<0,&-\infty<x_{1}<\infty,\\ (\rho_{+},\,{\boldsymbol{m}}_{+}),&x_{2}>0,&-\infty<x_{1}<\infty,\end{array}\right.$ (3.5)
where $\rho_{\pm},\,{\boldsymbol{m}}_{\pm}$ are constants.
Lemma 3.1.
$\tilde{{\boldsymbol{v}}}\otimes\tilde{{\boldsymbol{v}}}-\tilde{{\boldsymbol{u}}}<\frac{C}{2}Id.$ (3.6)
${\boldsymbol{v}}$ and ${\boldsymbol{u}}$ vanish identically outside $\Lambda$.
$(\tilde{{\boldsymbol{v}}}+{\boldsymbol{v}})\otimes(\tilde{{\boldsymbol{v}}}+{\boldsymbol{v}})-(\tilde{{\boldsymbol{u}}}+{\boldsymbol{u}})=\frac{C}{2}Id$  a.e. in $\Lambda$.
Several observations may be listed.
Remark 3.1 (Remarks on Lemma 3.1).
  1. 1.
    Lemma 3.1(i) states that the maps $({\boldsymbol{v}},\,{\boldsymbol{u}})$ have compact support $\Lambda$. In particular, these maps are restricted to a compact time interval and vanish smoothly outside this interval. Markfelder and Pellhammer [14] exploited the compactness when constructing a counterexample to the claim in [9]. Additional comment is provided in Section 5.
  2. 2.
    On appeal to [2, Sect. 6], the solutions $({\boldsymbol{v}},\,{\boldsymbol{u}})$ to the “pressureless” gas dynamics system of Lemma 3.1$(ii)$ may be taken as periodic in $x_{1}$. For later reference, we let the period be $2L_{3}$.
  3. 3.
    The reason for non-uniqueness of the convex integration solutions originates in the assertion in Lemma 3.1 that there are an infinite number of functions $({\boldsymbol{v}},\,{\boldsymbol{u}})$
  4. 4.
    The incompressible pressureless gas dynamics system specified in Lemma 3.1$(ii)$ is time reversible as proved on replacing $(t,\,x_{1},\,x_{2})$ by $(-t,\,-x_{1},\,-x_{2})$. Also, shocks cannot form in contrast to the usual time irreversible gas dynamics with non-zero pressure.
  5. 5.
The next task is to define the sub-solution of Chiodaroli and Kreml [2, Sect. 3]. Together with the corrugations of Lemma 3.1 it is an essential component of the convex integration process.
@@
where $\nu_{-}<\nu_{+}$ are arbitrary real numbers that correspond to the shock speeds and which depend upon $\bar{\rho}$. They are written $(\nu_{-}(\bar{\rho}),\,\nu_{+}(\bar{\rho}))$ to emphasize this dependence.
@@
where $\rho_{1},\,{\boldsymbol{v}}_{1},\,{\boldsymbol{u}}_{1}$ are constants with $\rho_{1}>0$ and where ${\boldsymbol{u}}_{-}={\boldsymbol{v}}_{-}\otimes{\boldsymbol{v}}_{-}-(1/2)|{\boldsymbol{v}}_{-}|^{2}{\boldsymbol{I}d}$, with a similar definition for ${\boldsymbol{u}}_{+}$.
@@
${\boldsymbol{v}}_{1}\otimes{\boldsymbol{v}}_{1}-{\boldsymbol{u}}_{1}<\frac{C}{2}{\boldsymbol{I}d}.$ (3.11)
@@
$\partial_{t}(\bar{\rho}\bar{{\boldsymbol{v}}})+div_{x}(\bar{\rho}\bar{{\boldsymbol{u}}})+\nabla_{x}\left(p(\bar{\rho})+\frac{1}{2}\left(C\rho_{1}{\boldsymbol{I}d}_{P_{1}}+\bar{\rho}|\bar{{\boldsymbol{v}}}|^{2}{\boldsymbol{I}d}_{P_{+}\cup P_{-}}\right)\right)$ $=$ $0$ (3.13)

### v4

@@ 1 Introduction
The paper [9] introduces a version of the least action principle as a selection criterion in initial value problems possessing non-unique solutions. Interest arises from examples that use convex integration techniques to obtain families of non-unique wild solutions to the Euler system (De Lellis and Székelyhidi [4, 5]), and non-unique solutions to the Navier-Stokes equations (Buckmaster and Vicol [1]). Convex integration is also employed by Chiodaroli and Kreml [2] to investigate the two-dimensional Riemann initial value problem for the isentropic compressible Euler system. They construct a large set of entropic wild solutions to the Riemann problem. Consequently, in these problems it is desirable to look for an admissibility criterion that distinguishes a particular solution from the many possible. Because of its mathematical simplicity, we are motivated to seek an admissibitlity criterion that selects the classical 2-shock solution.
Other admissibility criteria in fluid dynamics include the principles of minimum potential energy or entropy, various rate of change criteria, and that of least action (LAAP) developed in [9]. The latter principle applied to the initial value problem considered in [2] selects the 2-shock solution provided the solutions obtained by convex integration are the only other ones admitted for comparison. On the other hand, Chiodaroli and Kreml [2] show that there exist Riemann data for which there are wild solutions that possess an entropy rate lower than the 2-shock solution which therefore is not selected by Dafermos’s entropy rate criterion [3]. Nevertheless, LAAP selects the 2-shock in preference to wild solutions in particular cases.
LAAP, however, does not select the 2-shock solution when the set of comparison solutions is enlarged to include other weak solutions in addition to the 2-shock and convex integration solutions. Markfelder and Pellhammer [14] recently constructed hybrid solutions to the same Riemann initial value problem for which, subject to prescribed initial data, LAAP does not identify the 2-shock solutions as being uniquely admissible. In fact, this result indicates that LAAP does not provide an admissible solution to the Riemann problem. Accordingly, an alternative to LAAP must be sought when the 2-shock solution in this Riemann problem is preferred.
@@
Section 2 collects basic definitions, states the least action admissibility principle (LAAP) derived in [9] and formulates the new least action admissibility principle (LAAP0) and two closely related rate criteria. Section 3, after a brief description of convex integration, introduces the two-dimensional Riemann initial value problem for the barotropic and isentropic compressible Euler equations of gas dynamics and summarises further material from [2] including the construction of sub-solutions. The 2-shock solution is also described. Justification for the introduction of LAAP0 is provided in Section 4, while Section 5 defines the relevant notion of action and applies LAAP0 to the Riemann problem treated in [2] to establish the admissibility of the 2-shock solution.
Normally accepted notation is introduced without comment.
@@ 2 Least action admissibility criteria
$\frac{d{\boldsymbol{u}}}{dt}=F({\boldsymbol{u}}),\qquad{\boldsymbol{u}}(t_{0})={\boldsymbol{u}}_{0},\qquad t\in(t_{0},\,T],$ (2.1)
where $F({\boldsymbol{u}})$ is a given operator and $t_{0},\,T]$ is the maximal time interval of existence of solutions belonging to $S$. Global existence is not assumed so that possibly $T<\infty.$. Later, the system (2.1) is specialised to be the Euler-Lagrange equation derived from a least action principle.
Suppose that $S$ is not a singleton, which implies that (2.1) has non-unique solutions. The objective is to formulate admissibility criteria stated in terms of the action whose strict forms select unique elements of $S$. The definition of the action to be used is postponed to Section 5. For present purposes, it is sufficient to assume that for each solution ${\boldsymbol{u}}\in S$, the action $A({\boldsymbol{u}})(t_{0},\,t)$ vanishes at $t=t_{0}$ for all ${\boldsymbol{u}}\in S$.
To start, it is convenient to recall the Least Action Admissibility Principle (LAAP) introduced in [9] and applied to the Riemann initial value problem for compressible Euler systems and to Dafermos’ nonlinear oscillator [8]. For fixed final time $t_{1}$, we have:
@@
In a time interval $t_{0},t_{1}]$, where $t_{1}\in{\mathbb{R}}$ is less than or equal to the infimum of the maximal time of existence for solutions in $S$, a solution ${\boldsymbol{u}}\in S$ is LAAP-admissible if the action $A({\boldsymbol{u}})(t_{0},t_{1})$ is not greater than the action of all other solutions in $S$. The solution ${\boldsymbol{u}}\in S$ is strictly LAAP-admissible when the inequality is strict.
In the context of the wild solutions of the Riemann problem of [2], in a number of cases LAAP identifies the 2-shock solution as the only strictly LAAP-admissible solution for all $t_{1}>t_{0}$. But when the set $S$ includes solutions other than the wild solutions constructed in [2] and the 2-shock solution, LAAP, as explained in Section 1, can be ineffective. Obviously, the action depends upon the value chosen for $t_{1}$. In fact, as discussed in the next Section, this is the main feature in the construction of Markfelder and Pellhammer [14]: some choices of $t_{1}$ allow the one-dimensional 2-shock solution to be LAAP-admissible, while others do not.
Here we formulate a local in time version of LAAP, called LAAP0, for which the 2-shock solution is admissible in particular cases of the Riemann problem treated in [2]. It is equally admissible when the set $S$ is enlarged to include the hybrid solutions of [14].
To define LAAP0 we need the following definitions. (Note that in Definition 2.2 and subsequently the time $t_{1}({\boldsymbol{v}})$ depends upon the solution $\boldsymbol{v}$}, and $t_{0}$ is such that $t_{0},T]$ is the maximal existence interval.)
@@
2. For a class of solutions more general than considered here, a definition of entropy rate is presented in [7, Sects 1.2,1.3] when the time derivative only exists almost everywhere. The corresponding general definition of the action rate is obtained by replacing the energy $E$ by the action $A$ in the discussion of [7].
@@
We next apply the principles introduced in Definitions 2.2 and 2.3 to the two-dimensional Riemann problem for compressible isentropic Euler equation studied in e.g., [2, 9, 14].
3 The Riemann problem for a compressible barotropic Euler system
This Section defines the 2-dimensional Riemann problem and the isentropic form to be analysed. For convenience, several subsections are devoted to summarising various solutions and their relevant properties required in the later application of LAAP0. In particular, we consider solutions constructed by Chiodaroli and Kreml [2] using convex integration techniques, the two main elements of which are first briefly reviewed.
@@ 3.1 Convex integration
The two main elements of the convex integration procedure pertinent to the present study are the concepts of a sub-solution and of corrugation. For equation (2.1) a sub-solution is the function $\boldsymbol{w}$} that satisfies the inequality
@@
By corrugation we mean an iterative scheme of superposing on the sub-solution a sequence of spatio-temporal oscillations at decreasingly small scales such that in the limit the combined subsolution and oscillation converges to an appropriately defined weak (exact) solution to the equation of interest. In Nash’s fundamental paper [15] on isometric embedding of Riemannian manifolds, sub-solutions correspond to short embeddings. Nash produced a corrugation algorithm which he termed stages. Gromov [11, 12] generalised the technique, while De Lellis and Székelyhidi [4, 5] applied the general theory of convex integration to construct wild solutions to the Euler equations. The lecture notes by Markfelder [13] may be consulted for details. An overview is provided in the survey by De Lellis and Székelyhidi [6].
Chiodaroli and Kreml [2] apply the convex integration procedure to the 2-dimensional Riemann problem for the isentropic compressible Euler system in the case when it admits a 2-shock solution. We briefly explain their construction of sub-solutions and of weak solutions and refer to the latter as wild solutions.
@@ 3.2 Isentropic compressible Euler system
In what follows, the independent variables are time $t$ and position $\boldsymbol{x}$} such that $(t,{\boldsymbol{x}})\in[0,\infty)\times{\mathbb{R}}^{2}$. Let $\rho>0$ be the unknown mass density, $\boldsymbol{v}$} the fluid velocity, and $p(\rho)$ the constitutively defined pressure. In an obvious notation, the two dimensional compressible barotropic Euler system of gas dynamics is given by
$\partial_{t}\rho+div_{x}(\rho{\boldsymbol{v}})$ $=$ $0,$ (3.1)
$\partial_{t}(\rho{\boldsymbol{v}})+div_{x}\left(\rho{\boldsymbol{v}}\otimes{\boldsymbol{v}}\right)+\nabla_{x}[p(\rho)]$ $=$ $0,$ (3.2)
to which are adjoined the initial conditions
$(\rho,\,{\boldsymbol{v}})(0,\,\cdot)=(\rho_{0},\,{\boldsymbol{v}}_{0}).$ (3.3)
The pair $(\rho,\,{\boldsymbol{v}})$ is a weak solution to the initial value problem for the system (3.1)-(3.3) when it satisfies this system in the sense of distributions.
We consider entropic weak solutions to the initial value problem (3.1)-(3.3), that is, weak solutions that satisfy the energy-entropy condition in the sense of distributions (see for example [2, 3, 13, 14]):
@@
Chiodaroli and Kreml [2] study the isentropic form of the barotropic system (3.1) and (3.2) for which $p(\rho)=\rho^{\gamma}$, where the constant $\gamma$ satisfies $1<\gamma\leq 3$. The Riemann initial data is
$(\rho_{0},\,{{\boldsymbol{v}}_{0}})=\left\{\begin{array}[]{lll}(\rho_{-},\,{{\boldsymbol{v}}_{-}}),&x_{2}<0,&-\infty<x_{1}<\infty,\\ (\rho_{+},\,{{\boldsymbol{v}}_{+}}),&x_{2}>0,&-\infty<x_{1}<\infty,\end{array}\right$. (3.5)
where $\rho_{\pm},\,{{\boldsymbol{v}}_{\pm}}$ are constants.
In [2, (2.47), Lemma 2.4(5)], Chiodaroli and Kreml show that for any Riemann data satisfying the necessary condition
$(v_{-}-v_{+})^{2}\rho_{+}\rho_{-}-(\rho_{+}-\rho_{-})(p(\rho_{+})-p(\rho_{-}))>0,$ (3.6)
there exists a unique self-similar solution to the initial value problem, the 2-shock solution, given by
$\left\{\begin{array}[]{lllll}\rho=\rho_{-},&v_{b}=v_{{-}},&t>0,&-\infty<x_{1}<\infty,&x_{2}<\nu_{-}t,\\ \rho=\rho_{+},&v_{b}=v_{{+}},&t>0,&-\infty<x_{1}<\infty,&x_{2}>\nu_{+}t,\\ \rho=\rho_{m},&v_{b}=v_{m},&t>0,&-\infty<x_{1}<\infty,&\nu_{-}t<x_{2}<\nu_{+}t.\end{array}\right$.
Here ${\boldsymbol{v}}=(0,\,v_{b})$; $\rho_{m}$ and $v_{m}$ are found from [2, (2.48), (2.49)] and the speeds of the shocks $\nu_{\pm}$ are determined from the Riemann data by the Rankine-Hugoniot conditions; see [2, (2.29), (2.32)].
3.3 Fan sub-solutions
The next task is to define the sub-solution of Chiodaroli and Kreml [2, Sect. 3]. Together with the corrugations of Lemma 3.1 below it is an essential component of obtaining weak solutions by the convex integration process.
@@ 3.3 Fan sub-solutions
where $\nu_{-}<\nu_{+}$ are arbitrary real numbers that correspond to the shock speeds.
@@
where $\rho_{1},\,{\boldsymbol{v}}_{1},\,{\boldsymbol{u}}_{1}$ are constants with $\rho_{1}>0$ and where ${\boldsymbol{u}}_{-}={\boldsymbol{v}}_{-}\otimes{\boldsymbol{v}}_{-}-(1/2)|{\boldsymbol{v}}_{-}|^{2}\mathrm{Id}$, with a similar definition for $\boldsymbol{u}$$_{+}$.
@@
${\boldsymbol{v}}_{1}\otimes{\boldsymbol{v}}_{1}-{\boldsymbol{u}}_{1}<\frac{C}{2}\mathrm{Id}.$ (3.11)
@@
$\partial_{t}(\bar{\rho}\bar{{\boldsymbol{v}}})+div_{x}(\bar{\rho}\bar{{\boldsymbol{u}}})+\nabla_{x}\left(p(\bar{\rho})+\frac{1}{2}\left(C\rho_{1}\textup{I}_{P_{1}}+\bar{\rho}|\bar{{\boldsymbol{v}}}|^{2}\textup{I}_{P_{+}\cup P_{-}}\right)\right)$ $=$ $0.$ (3.13)
@@
3.4 Corrugations
In their article Chiodaroli and Kreml [2] prove non-uniqueness of weak solutions to the Riemann initial value problem with initial conditions (3.5), i.e. not only is there the classical 2-shock solution given in Section 3.2, but also an infinite number of weak solutions that possess non-trivial $x_{1}$ dependence. Their proof, based upon convex integration, consists of two pieces: (i) a base sub-solution and (ii) an infinite sequence of corrugations which when combined with (i) yields an infinite number of solutions to the initial value problem. Here we recall these two concepts within the context of our Riemann initial value problem for the isentropic Euler equations. We emphasize that the corrugations combined with even a single sub-solution yield an infinite number of weak solutions to initial value problem with initial conditions (3.5).
Lemma 3.1.
$\tilde{{\boldsymbol{v}}}\otimes\tilde{{\boldsymbol{v}}}-\tilde{{\boldsymbol{u}}}<\frac{C}{2}\mathrm{Id}.$ (3.15)
$\boldsymbol{v}$} and $\boldsymbol{u}$} vanish identically outside $\Lambda$.
$(\tilde{{\boldsymbol{v}}}+{\boldsymbol{v}})\otimes(\tilde{{\boldsymbol{v}}}+{\boldsymbol{v}})-(\tilde{{\boldsymbol{u}}}+{\boldsymbol{u}})=\frac{C}{2}\mathrm{Id}$  a.e. in $\Lambda$.
Several observations may be listed.
Remark 3.1 (Remarks on Lemma 3.1).
  1. 1.
    Markfelder and Pellhammer [14] exploited in their discussion that $({\boldsymbol{v}},\,{\boldsymbol{u}})$ of Lemma 3.1 vanish after a finite time, allowing them to construct a hybrid weak solution. Additional comment is provided in Section 5.
  2. 2.
    On appeal to [2, Sect. 6], the solutions $({\boldsymbol{v}},\,{\boldsymbol{u}})$ to the system of Lemma 3.1 may be taken as periodic in $x_{1}$. For later reference, we let the period be $2L_{3}$.
  3. 3.
    The reason for non-uniqueness of the convex integration solutions originates in the assertion in Lemma 3.1 that there are an infinite number of functions $({\boldsymbol{v}},\,{\boldsymbol{u}})$ .
  4. 4.
    The system specified in Lemma 3.1 is time reversible as proved on replacing $(t,\,x_{1},\,x_{2})$ by $(-t,\,-x_{1},\,-x_{2})$.
  5. 5.
3.5 Convex integration solutions